# One Subgoal at a Time: Zero-Shot Generalization to Arbitrary Linear Temporal Logic Requirements in Multi-Task Reinforcement Learning

### v1

@@ Abstract
Generalizing to complex and temporally extended task objectives and safety constraints remains a critical challenge in reinforcement learning (RL). Linear temporal logic (LTL) offers a unified formalism to specify such requirements, yet existing methods are limited in their abilities to handle nested long-horizon tasks and safety constraints, and cannot identify situations when a subgoal is not satisfiable and an alternative should be sought. In this paper, we introduce GenZ-LTL, a method that enables zero-shot generalization to arbitrary LTL specifications. GenZ-LTL leverages the structure of Büchi automata to decompose an LTL task specification into sequences of reach-avoid subgoals. Contrary to the current state-of-the-art method that conditions on subgoal sequences, we show that it is more effective to achieve zero-shot generalization by solving these reach-avoid problems one subgoal at a time through proper safe RL formulations. In addition, we introduce a novel subgoal-induced observation reduction technique that can mitigate the exponential complexity of subgoal-state combinations under realistic assumptions. Empirical results show that GenZ-LTL substantially outperforms existing methods in zero-shot generalization to unseen LTL specifications.
@@ 1 Introduction
In this paper, we propose a novel framework for learning RL policies that generalize in a zero-shot manner to satisfy arbitrary LTL specifications, including both finite-horizon and infinite-horizon tasks. The core idea is to exploit the structure of the equivalent Büchi automata to decompose an LTL specification into individual subgoals, each comprising a reach component that defines task progression and an avoid component that encodes safety constraints. A goal-conditioned policy is then trained to achieve the subgoals one at a time through a safe RL formulation that can effectively handle the safety constraints and optimize for efficient satisfaction of the subgoal. During training, we sample different subgoals from the space of possible subgoals to train the goal-conditioned policy. In addition, we introduce a novel subgoal-induced observation reduction technique to address the exponential complexity of observation-subgoal combinations. This technique exploits symmetry in the state observations and extracts information only relevant to the current subgoals. At test time, given an arbitrary LTL specification, we use its equivalent Büchi automaton as well as the learned value function to guide the subgoal selection. Furthermore, we introduce a timeout-based subgoal-switching mechanism that tracks subgoal satisfaction and enables the switching to alternative subgoals if the current one turns out to be unsatisfiable, a situation that existing methods fail to manage. Our main contributions are summarized as follows.
@@ 2 Related Work
Specification-guided and goal-conditioned RL. Various methods have been proposed to guide RL agents by extending traditional goal-conditioned RL (liu2022goal, ) to temporal-logic goals. Early works focused on learning policies for a single and fixed specification, including methods based on LTL specifications and automaton-like models (hasanbeig2018LogicallyConstrained, ; hahn2023Mungojerrie, ; voloshin2023eventual, ; shao2023Sample, ; shah2024LTLConstrained, ; bagatella2024Directed, ), and methods leveraging quantitative semantics of specifications (li2017reinforcement, ; jothimurugan2019composable, ; shah2024deep, ; guo2024temporal, ). More recently, the importance of adapting to changing conditions in real-world applications has brought increased attention to zero-shot generalization to new specifications. Existing approaches can be broadly categorized into subgoal decomposition (leon2021Systematic, ; leon2022Nutshell, ; liu2024Skill, ; qiu2023instructing, ; jackermeierdeepltl, ) and direct specification/automata encoding (vaezipoor2021ltl2action, ; yalcinkaya2024compositional, ; yalcinkaya2025provably, ). Our work follows the direction of subgoal decomposition, but shows that it is more effective to solve the subgoals one at a time through proper safe RL formulations and subgoal-induced observation reduction.
RL with safety constraints. Various methods have been proposed to solve RL problems with safety constraints, such as policy optimization-based methods zhang2020first ; Yang2020Projection-Based ; jayant2022model ; xiong2023provably , Gaussian processes-based approaches (wachi2018safe, ; zhao2023probabilistic, ), and control theory-based methods, including Hamilton-Jacobian (HJ) reachability fisac2019bridging ; yu2022reachability ; so2023solving ; ganai2023iterative ; zheng2024safe and control barrier functions (gangopadhyay2022safe, ; cheng2023safe, ). We point the readers to gu2022review for a comprehensive survey of safe RL techniques. In the context of formal specification-guided RL, the interplay between task progression and safety constraints remains underexplored. Most existing methods simply assign a positive reward when a specification or a subgoal is satisfied, and a negative reward at episode termination if unsatisfied. This approach is ineffective in achieving safety (achiam2017constrained, ; dalal2018safe, ; bharadhwaj2020conservative, ), especially when the task progression objective and the safety requirements exert conflicting influences on policy learning. In this paper, we show how to prioritize safety satisfaction properly by incorporating Hamilton-Jacobi reachability constraints.
@@ 4.1 Reach-Avoid Subgoal Construction from Büchi Automaton
[…]
@@ 4.2 Subgoal-Induced Observation Reduction
[…]
@@
In this paper, we introduce a novel subgoal-induced observation reduction technique that can circumvent this exponential complexity. The key observation is that, ultimately, each subgoal reduces to a reach-avoid problem, where there is some state that we want to reach while avoiding some other states. In other words, what really matters is whether a state is ‘reach’ or ‘avoid’, and not its label $L(s)$ once a subgoal is fixed. Thus, we can perform a symbolic simplification of the state/observation space along with the subgoal while preserving information essential for solving the reach-avoid problem. Formally, for an MDP state $s$, we assume that it can be partition into a proposition-independent component $s_{\neg AP}$ and a proposition-dependent component $s_{AP}$ (e.g, the blue and green sensors). For the proposition-dependent component, we further assume that it can be broken down into individual observations each corresponding to an element in $2^{AP}$, and the observations are produced by identical observation functions under output transformation. We say that two observation functions (with a slight abuse of notation of $s$), $o_{1}(s)\in\mathbb{R}^{k}$ and $o_{2}(s)\in\mathbb{R}^{k}$ are identical under an output transformation $g$ if $\forall s,o_{1}(s)=g(o_{2}(s))$. For instance, suppose $o_{1}(s)\in\mathbb{R}^{360}$ is a LiDAR sensor for sensing the shortest distance to a blue state along each of its 360360360 beams, and $o_{2}(s)\in\mathbb{R}^{360}$ is the same type of LiDAR sensor for sensing the shortest distance to a green state but rotated $180^{\circ}$, then $g$ is the $180^{\circ}$ rotation (a permutation of the beam indices). Then, we can reduce the proposition-dependent component (of size $k\times 2^{|AP|}$) and the encoding of a reach-avoid subgoal (of size ($|AP|+2^{|AP|}$) into just two observations each of size $k$, where the first observation is an observation corresponding to the reach subgoal and the second observation is a fused observation of the observations corresponding to the avoid subgoals. For simplicity of deriving a fusion operator that preserves task-relevant semantics such as distances to avoid zones, we assume that the $o(s)$ are coordinate-wise monotonic. The actual fusion operator $f_{\text{}}$ is environment/sensor-specific – in the LiDAR example, it will be an elementwise $\min$ over $\mathbb{R}^{k}$. Implementation details are provided in Appendix B. Empirically, we demonstrate the benefits of this reduction, when it is applicable, in Sections 5.3 and Apendix C.
@@ 4.3 Goal-Conditioned Policy Learning with Reachability Constraints
[…]
@@
[…]
@@
where $A_{r}^{\pi}(s^{\sigma},a):=Q_{r}^{\pi}(s^{\sigma},a)-V_{r}^{\pi}(s^{\sigma})$ and $A_{h}^{\pi}(s^{\sigma},a):=Q_{h}^{\pi}(s^{\sigma},a)-V_{h}^{\pi}(s^{\sigma})$ are the advantages functions with $Q_{h}^{\pi}(s^{\sigma},a):=\max_{t\in\mathbb{N}}h(s^{\sigma}_{t}),s_{0}=s^{\sigma},a_{0}=a,a_{t}\sim\pi$ and $J_{h}(\pi):=\max_{t\in\mathbb{N}}h(\cdot),a\sim\pi$ is the maximum safety violation for a trajectory under policy $\pi$. By maximizing the discounted cumulative reward, the policy is optimized to effectively satisfy the reach subgoal, while the safety constraints ensures that the avoid subgoals are satisfied. Then we incorporate the clipped surrogate objective schulman2017proximal to handle the trust-region constraint and a state-dependent Lagrangian multiplier yu2022reachability to enforce the state-wise safety constraint:
[…]
@@
Subgoal switching. In general, we cannot determine a priori whether a selected subgoal is satisfiable in the environment. To cope with this issue, we introduce a mechanism to identify unsatisfiable subgoals and allow the agent to switch to alternative ones. The mechanism is triggered if the current subgoal is not satisfied within a timeout threshold, which is set to $(1+\epsilon_{\text{scale}})\cdot\mu_{\text{subgoal}}$, where $\mu_{\text{subgoal}}$ denotes the maximum number of steps required to complete a subgoal after policy convergence during training, and $\epsilon_{\text{scale}}$ is a scaling factor. When the current reach subgoal times out, the best subgoal from the remaining reach subgoals in $A^{+}$ is selected. This procedure is sound because the avoid subgoals are respected during the attempt for the prior reach subgoal(s). Appendix B.2 provides further implementation details.
@@ 5 Experiments
Environments. Our environments include the LetterWorld environment vaezipoor2021ltl2action , a $\times 7$ 7 discrete grid world with discrete actions where letters are placed at randomly sampled positions, and the ZoneEnv environment vaezipoor2021ltl2action , a high-dimensional environment with lidar observations, where a robotic agent with a continuous action space must navigate between randomly positioned colored regions. We exclude the FlatWorld environment jackermeierdeepltl ; voloshin2023eventual from our experiments, as its configuration is fixed and the state space includes only the agent’s $(x,y)$-position, without any information about the environment state. We also consider a variant of ZoneEnv that contains overlapping regions which is by design more challenging than FlatWorld. Additional details are provided in Appendix A.1.
@@
LTL specifications. We consider a wide variety of LTL specifications with varying complexity, including (1) sequential reach-only specifications of the form $\mathsf{F}\,(\alpha_{1}\land(\mathsf{F}\,\alpha_{2}\land(\ldots\land\mathsf{F}\,\alpha_{n}))$; (2) sequential reach-avoid specifications of the form $\neg\alpha_{1}\;\mathsf{U}\;(\alpha_{2}\land(\neg\alpha_{3}\;\mathsf{U}\;(\alpha_{4}\land\ldots\land(\neg\alpha_{n-1}\;\mathsf{U}\;\alpha_{n}))$, and (3) complex specifications that combine reach-only and reach-avoid properties. We also conduct a separate evaluation on infinite-horizon tasks that span a diverse set of LTL objectives, including safety, liveness, and their combinations. Details of the exact specifications we use in each environment are given in Appendix A.2, and more visualization of trajectories are provided in Appendix D. In this section, all specifications are satisfiable.

### v6

@@ Abstract
Generalizing to complex and temporally extended task objectives and safety constraints remains a critical challenge in reinforcement learning (RL). Linear temporal logic (LTL) offers a unified formalism to specify such requirements, yet existing methods are limited in their abilities to handle nested long-horizon tasks and safety constraints, and cannot identify situations when a subgoal is not satisfiable and an alternative should be sought. In this paper, we introduce GenZ-LTL, a method that enables zero-shot generalization to arbitrary LTL specifications. GenZ-LTL leverages the structure of Büchi automata to decompose an LTL task specification into sequences of reach-avoid subgoals. Contrary to the current state-of-the-art method that conditions on subgoal sequences, we show that it is more effective to achieve zero-shot generalization by solving these reach-avoid problems one subgoal at a time through proper safe RL formulations. In addition, we introduce a novel subgoal-induced observation reduction technique that can mitigate the exponential complexity of subgoal-state combinations under realistic assumptions. Empirical results show that GenZ-LTL substantially outperforms existing methods in zero-shot generalization to unseen LTL specifications. The code is available at https://github.com/BU-DEPEND-Lab/GenZ-LTL.
@@ 1 Introduction
In this paper, we propose a novel framework, called GenZ-LTL, for learning RL policies that generalize in a zero-shot manner to satisfy arbitrary LTL specifications, including both finite-horizon and infinite-horizon tasks. We adopt the subgoal decomposition approach, which exploits the structure of the equivalent Büchi automaton to decompose an LTL specification into individual subgoals, each comprising a reach component that defines task progression and an avoid component that encodes safety constraints. We show that training a goal-conditioned policy to achieve the subgoals one at a time, which is often perceived as myopic by existing literature, is more effective in general than conditioning the policy on subgoal sequences jackermeierdeepltl or the entire automaton yalcinkaya2024compositional. This somewhat surprising result stems from two key innovations in this paper: a proper safe RL formulation with state-wise constraints for solving the individual reach-avoid subgoals, and a novel subgoal-induced observation reduction technique to address the exponential complexity of observation-subgoal combinations. The former builds on recent advances of Hamilton-Jacobi reachability in RL ganai2024hamilton whereas the latter exploits symmetry in the state observations to extract information only relevant to the immediate subgoal. To zero-shot generalize to arbitrary LTL specifications at test time, we use the equivalent Büchi automata as well as the learned value functions to guide the subgoal selection. Our main contributions are summarized as follows.
@@ 2 Related Work
Specification-guided and goal-conditioned RL. Various methods have been proposed to guide RL agents by extending traditional goal-conditioned RL (liu2022goal) to temporal-logic goals. Early works focused on learning policies for a single and fixed specification, including methods based on LTL specifications and automaton-like models (hasanbeig2018LogicallyConstrained; hahn2023Mungojerrie; voloshin2023eventual; shao2023Sample; shah2024LTLConstrained; bagatella2024Directed), and methods leveraging quantitative semantics of specifications (li2017reinforcement; jothimurugan2019composable; shah2024deep; guo2024temporal). More recently, the importance of adapting to changing conditions in real-world applications has brought increased attention to zero-shot generalization to new specifications. Existing approaches can be broadly categorized into subgoal decomposition (leon2021Systematic; leon2022Nutshell; liu2024Skill; qiu2023instructing; jackermeierdeepltl) and direct specification/automata encoding (vaezipoor2021ltl2action; yalcinkaya2024compositional; yalcinkaya2025provably). Our work follows the direction of subgoal decomposition, but shows that it is more effective to solve the subgoals one at a time through proper safe RL formulations and subgoal-induced observation reduction.
RL with safety constraints. Various methods have been proposed to solve RL problems with safety constraints, such as policy optimization-based methods zhang2020first; Yang2020Projection-Based; jayant2022model; xiong2023provably, Gaussian processes-based approaches (wachi2018safe; zhao2023probabilistic), and control theory-based methods, including Hamilton-Jacobian (HJ) reachability fisac2019bridging; yu2022reachability; so2023solving; ganai2023iterative; zheng2024safe and control barrier functions (gangopadhyay2022safe; cheng2023safe). We point the readers to gu2022review for a comprehensive survey of safe RL techniques. In the context of formal specification-guided RL, the interplay between task progression and safety constraints remains underexplored. Most existing methods simply assign a positive reward when a specification or a subgoal is satisfied, and a negative reward at episode termination if unsatisfied. This approach is ineffective in achieving safety (achiam2017constrained; dalal2018safe; bharadhwaj2020conservative), especially when the task progression objective and the safety requirements exert conflicting influences on policy learning. In this paper, we show how to prioritize safety satisfaction properly by incorporating Hamilton-Jacobi reachability constraints.
@@ 4.1 Reach-Avoid Subgoal Construction from Büchi Automaton
Büchi automata can be represented as directed state-transition graphs qiu2023instructing, and an accepted run can be viewed as having a lasso structure with a cycle containing an accepting state and a prefix path that leads to the cycle. Similar to existing approaches qiu2023instructing; jackermeierdeepltl, we use depth-first search (DFS) over the state-transition graph to enumerate all possible lasso structures starting from a given state $q$. From these lasso paths, we extract the set of reach subgoals, denoted by $A^{+}$, following the definition in Section 3. Similarly, we extract the set of avoid subgoals $A^{-}$ by considering transitions to a state that is different from $q$ and not along any of the lasso paths. Each pair $(\alpha^{+}\in A^{+},A^{-})$ then becomes a candidate reach-avoid subgoal. To satisfy a subgoal $(\alpha^{+},A^{-})$, the agent must visit a state $s$ such that $L(s)=\alpha^{+}$, while avoiding any states labeled with propositions in $A^{-}$. For example, for $(A^{+}=\{\{a,b\},\{c\}\},A^{-}=\{\{d\},\{e\}\})$, it forms subgoals $\sigma_{1}=(\{a\land b\},A^{-})$ and $\sigma_{2}=(\{c\},A^{-})$, and the agent can choose to reach a state labeled with $a\wedge b$ or a state labeled with $c$, while avoiding states labeled with $d$ and $e$ in both cases. While our method follows a similar subgoal decomposition approach as prior works, the way that we solve the reach-avoid subgoals is significantly different. In fact, a straightforward way is to encode a reach-avoid subgoal as a bitvector of dimension $|AP|+2^{|AP|}$, where the first $|AP|$ bits encodes the reach subgoal, and the rest of the bits encode the avoid subgoal.
@@ 4.2 Subgoal-Induced Observation Reduction
Given a reach-avoid subgoal, our next step is to train a subgoal-conditioned policy based on the MDP state or observation of the MDP state, and the subgoal. However, notice that the number of possible subgoals grows exponentially with the number of atomic propositions. Additionally, in order for the agent to complete a task in such a complex environment, it needs to be able to make observation of every state-subgoal combination, i.e. being able to determine the label $L(s)$ of a state $s$ that it observes (possibly indirectly through an observation function $o$). Otherwise, task completion cannot be determined in general. For instance, if the task requires the agent to reach a ‘green’ region, then an agent equipped with only ‘blue’ sensors cannot learn to accomplish this task if the green and blue regions are placed randomly in the environment. Thus, the number of observations that the policy needs to condition on also grows exponentially with $|AP|$, and directly learning a goal-conditioned policy this way quickly becomes intractable as $|AP|$ grows. While various dimensionality reduction techniques have been proposed in the RL literature allen2021learning; hansen2022bisimulation; zang2023understanding, they are based on abstraction or bisimulation of the MDP states, and do not consider state-subgoal combinations.
@@
In this paper, we introduce a novel subgoal-induced observation reduction technique that can circumvent this exponential complexity. The key observation is that, ultimately, each subgoal reduces to a reach-avoid problem, where there is some state that we want to reach while avoiding some other states. In other words, what really matters is whether a state is ‘reach’ or ‘avoid’, and not its label $L(s)$ once a subgoal is fixed. Thus, we can perform a symbolic simplification of the state/observation space along with the subgoal while preserving information essential for solving the reach-avoid problem. Formally, for an MDP state $s$, we assume that it can be partition into a proposition-independent component $s_{\neg AP}$ and a proposition-dependent component $s_{AP}$ (e.g, the blue and green sensors). For the proposition-dependent component, we further assume that it can be broken down into individual observations each corresponding to an element in $2^{AP}$, and the observations are produced by identical observation functions under output transformation. We say that two observation functions (with a slight abuse of notation of $s$), $o_{1}(s)\in\mathbb{R}^{k}$ and $o_{2}(s)\in\mathbb{R}^{k}$ are identical under an output transformation $g$ if $\forall s,o_{1}(s)=g(o_{2}(s))$. For instance, suppose $o_{1}(s)\in\mathbb{R}^{360}$ is a LiDAR sensor for sensing the shortest distance to a blue state along each of its $360$ beams, and $o_{2}(s)\in\mathbb{R}^{360}$ is the same type of LiDAR sensor for sensing the shortest distance to a green state but rotated $180^{\circ}$, then $g$ is the $180^{\circ}$ rotation (a permutation of the beam indices). Then, we can reduce the proposition-dependent component (of size $k\times 2^{|AP|}$) and the encoding of a reach-avoid subgoal (of size ($|AP|+2^{|AP|}$) into just two observations each of size $k$, where the first observation is an observation corresponding to the reach subgoal and the second observation is a fused observation of the observations corresponding to the avoid subgoals. For simplicity of deriving a fusion operator that preserves task-relevant semantics such as distances to avoid zones, we assume that the $o(s)$ are coordinate-wise monotonic. The actual fusion operator $f_{\text{}}$ is environment/sensor-specific – in the LiDAR example, as shown in Figure 2, it will be an elementwise $\min$ over $\mathbb{R}^{k}$. Implementation details are provided in Appendix B. Empirically, we demonstrate the benefits of this reduction, when it is applicable, in Sections 5.3 and Apendix C.
Figure 2: Illustration of the subgoal-induced observation reduction for LiDAR observations. The environment provides the raw state consisting of a proposition-independent component $s_{\neg AP}$ and a proposition-dependent component $s_{AP}$. Given a subgoal $\sigma$, $s_{AP}$ is reduced to two vectors: a reach observation and a fused avoid observation. Together with the $s_{\neg AP}$, these form the agent’s input.
@@ 4.3 Goal-Conditioned Policy Learning with Reachability Constraints
$\text{ s.t. }V_{h}^{\pi}(s^{\sigma})\leq 0,\forall s^{\sigma}\in\mathcal{S}_{f}^{\sigma}\cap\mathcal{S}_{0}^{\sigma},\forall\sigma\sim\mathrm{Unif}(\xi)$
@@
$\pi_{k+1}=\underset{\pi}{\arg\max}\ \mathbb{E}_{\sigma\sim\mathrm{Unif}(\xi),s\sim d^{\pi_{k}},a\sim\pi_{k}}\left[\frac{\pi}{\pi_{k}}A_{r}^{\pi_{k}}(s^{\sigma},a)\right]$ (2)
$\text{s.t.}\quad\mathbb{E}_{\sigma\sim\mathrm{Unif}(\xi),s\sim d^{\pi_{k}}}[\mathcal{D}_{KL}(\pi,\pi_{k})]\leq\epsilon,$
@@
where $A_{r}^{\pi}(s^{\sigma},a):=Q_{r}^{\pi}(s^{\sigma},a)-V_{r}^{\pi}(s^{\sigma})$ and $A_{h}^{\pi}(s^{\sigma},a):=Q_{h}^{\pi}(s^{\sigma},a)-V_{h}^{\pi}(s^{\sigma})$ are the advantages functions with $Q_{h}^{\pi}(s^{\sigma},a):=\max_{t\in\mathbb{N}}h(s^{\sigma}_{t}),s_{0}=s^{\sigma},a_{0}=a,a_{t}\sim\pi$ and $J_{h}(\pi):=\max_{t\in\mathbb{N}}h(\cdot),a\sim\pi$ is the maximum safety violation for a trajectory under policy $\pi$. By maximizing the discounted cumulative reward, the policy is optimized to effectively satisfy the reach subgoal, while the safety constraints ensures that the avoid subgoals are satisfied, Then we incorporate the clipped surrogate objective schulman2017proximal to handle the trust-region constraint and a state-dependent Lagrangian multiplier yu2022reachability to enforce the state-wise safety constraint:
$\underset{\lambda\geq 0}{\min}\ \underset{\pi}{\max}\ \mathbb{E}_{\sigma\sim\mathrm{Unif}(\xi),s\sim d^{\pi_{k}},a\sim\pi}\Big[$ $\min(\frac{\pi}{\pi_{k}}A_{r}^{\pi_{k}}(s^{\sigma},a),\ \text{clip}(\frac{\pi}{\pi_{k}},1-\epsilon,1+\epsilon)A_{r}^{\pi_{k}}(s^{\sigma},a))$ (3)
@@
Subgoal switching. In general, we cannot determine a priori whether a selected subgoal is satisfiable in the environment. To handle this issue, we introduce a mechanism to identify unsatisfiable subgoals and allow the agent to switch to alternative ones. The mechanism is triggered if the current subgoal is not satisfied within a timeout threshold, which is set to $(1+\epsilon_{\text{scale}})\cdot\mu_{\text{subgoal}}$, where $\mu_{\text{subgoal}}$ denotes the maximum number of steps required to complete a subgoal after policy convergence during training, and $\epsilon_{\text{scale}}$ is a scaling factor. When a reach subgoal times out, the best option from the remaining reach subgoals in $A^{+}$ is selected. This procedure is sound because the avoid subgoals are respected during the attempt for the prior reach subgoal(s). Appendix B.2 provides further details.
@@ 5 Experiments
Environments. Our environments include the LetterWorld environment vaezipoor2021ltl2action, a $7\times 7$ discrete grid world with discrete actions where letters are placed at randomly sampled positions, and the ZoneEnv environment vaezipoor2021ltl2action, a high-dimensional environment with lidar observations, where a robotic agent with a continuous action space must navigate between randomly positioned colored regions. We exclude the FlatWorld environment jackermeierdeepltl; voloshin2023eventual from our experiments, as its configuration is fixed and the state space includes only the agent’s $(x,y)$-position, without any information about the environment state. We also consider a variant of ZoneEnv that contains overlapping regions which is by design more challenging than FlatWorld. Additional details are provided in Appendix A.1.
@@
LTL specifications. We consider a wide variety of LTL specifications with varying complexity, including (1) sequential reach-only specifications of the form $\mathsf{F}\,(\alpha_{1}\land(\mathsf{F}\,\alpha_{2}\land(\ldots\land\mathsf{F}\,\alpha_{n}))$; (2) sequential reach-avoid specifications of the form $\neg\alpha_{1}\;\mathsf{U}\;(\alpha_{2}\land(\neg\alpha_{3}\;\mathsf{U}\;(\alpha_{4}\land\ldots\land(\neg\alpha_{n-1}\;\mathsf{U}\;\alpha_{n}))$, and (3) complex specifications that combine reach-only and reach-avoid properties. We also conduct a separate evaluation on infinite-horizon tasks that span a diverse set of LTL objectives, including safety, liveness, and their combinations. The specifications used for evaluation are listed in Table 1 and a full set of specifications is provided in Appendix A.2, with trajectory visualizations in Appendix D. In this section, all specifications are satisfiable.
Table 1: LTL specifications used for evaluation. For specifications in ZoneEnv, $\mathsf{b}$, $\mathsf{g}$, $\mathsf{m}$, and $\mathsf{y}$ denote $\mathsf{blue}$, $\mathsf{green}$, $\mathsf{magenta}$, and $\mathsf{yellow}$, respectively. Results are shown in Table 2 and 3.
  LetterWorld ZoneEnv
Finite-horizon $\varphi_{1}$ $\mathsf{F}\,\left(\mathsf{a}\land(\neg\mathsf{b}\;\mathsf{U}\;\mathsf{c})\right)\land\mathsf{F}\,\mathsf{d}$ $\varphi_{9}$ $(\mathsf{F}\,\mathsf{b})\land\left(\neg\mathsf{b}\;\mathsf{U}\;(\mathsf{g}\land\mathsf{F}\,\mathsf{y})\right)$
$\varphi_{2}$ $(\mathsf{F}\,\mathsf{d})\land\left(\neg\mathsf{f}\;\mathsf{U}\;(\mathsf{d}\land\mathsf{F}\,\mathsf{b})\right)$ $\varphi_{10}$ $\neg(\mathsf{m}\lor\mathsf{y})\ \;\mathsf{U}\;\ \left(\mathsf{b}\land\mathsf{F}\,\mathsf{g}\right)$
$\varphi_{3}$ $\neg\mathsf{a}\ \;\mathsf{U}\;\ \left(\mathsf{b}\land\left(\neg\mathsf{c}\;\mathsf{U}\;\left(\mathsf{d}\land(\neg\mathsf{e}\;\mathsf{U}\;\mathsf{f})\right)\right)\right)$ $\varphi_{11}$ $\neg\mathsf{g}\ \;\mathsf{U}\;\ \left((\mathsf{b}\lor\mathsf{m})\land(\neg\mathsf{g}\;\mathsf{U}\;\mathsf{y})\right)$
$\varphi_{4}$ $\left(\mathsf{a}\lor\mathsf{b}\lor\mathsf{c}\lor\mathsf{d}\Rightarrow\mathsf{F}\,\left(\mathsf{e}\land\mathsf{F}\,\left(\mathsf{f}\land\mathsf{F}\,\mathsf{g}\right)\right)\right)\ \;\mathsf{U}\;\ \left(\mathsf{h}\land\mathsf{F}\,\mathsf{i}\right)$ $\varphi_{12}$ $\left(\mathsf{g}\lor\mathsf{b}\Rightarrow\left(\neg\mathsf{y}\;\mathsf{U}\;\mathsf{m}\right)\right)\ \;\mathsf{U}\;\ \mathsf{y}$
$\varphi_{5}$ $\mathsf{F}\,\left(\mathsf{d}\land\left(\neg(\mathsf{a}\lor\mathsf{b})\;\mathsf{U}\;(\mathsf{b}\land(\neg\mathsf{e}\;\mathsf{U}\;\mathsf{c}))\right)\right)\land\mathsf{F}\,\left(\neg(\mathsf{f}\lor\mathsf{g}\lor\mathsf{h})\;\mathsf{U}\;\mathsf{a}\right)$ $\varphi_{13}$ $\mathsf{F}\,\left(\mathsf{g}\land\left(\neg(\mathsf{b}\lor\mathsf{y})\;\mathsf{U}\;(\mathsf{y}\land(\neg\mathsf{m}\;\mathsf{U}\;\mathsf{b}))\right)\right)\land\mathsf{F}\,\left(\neg\mathsf{g}\;\mathsf{U}\;\mathsf{y}\right)$
$\varphi_{6}$ $\mathsf{F}\,\left(\left(\mathsf{k}\land\left(\left(\neg\mathsf{b}\lor\mathsf{c}\right)\;\mathsf{U}\;\mathsf{f}\right)\right)\land\left(\neg(\mathsf{a}\lor\mathsf{e}\lor\mathsf{h})\;\mathsf{U}\;\mathsf{g}\right)\right)\land\mathsf{F}\,\mathsf{d}$ $\varphi_{14}$ $\mathsf{F}\,\left((\mathsf{b}\lor\mathsf{g})\land\left(\neg\mathsf{y}\;\mathsf{U}\;(\mathsf{b}\land(\neg(\mathsf{g}\lor\mathsf{m})\;\mathsf{U}\;\mathsf{m}))\right)\right)\land\mathsf{F}\,\left(\mathsf{y}\land(\neg\mathsf{b}\;\mathsf{U}\;\mathsf{g})\right)$
$\varphi_{7}$ $\neg(\mathsf{j}\lor\mathsf{b}\lor\mathsf{d})\ \;\mathsf{U}\;\ \left(\mathsf{a}\land(\neg\mathsf{c}\;\mathsf{U}\;(\mathsf{f}\land\mathsf{F}\,(\mathsf{g}\land(\neg\mathsf{d}\;\mathsf{U}\;\mathsf{e}))))\right)$ $\varphi_{15}$ $\neg(\mathsf{m}\lor\mathsf{y})\ \;\mathsf{U}\;\ \left(\mathsf{b}\land(\neg\mathsf{g}\;\mathsf{U}\;(\mathsf{y}\land\mathsf{F}\,(\mathsf{g}\land(\neg\mathsf{b}\;\mathsf{U}\;\mathsf{m}))))\right)$
$\varphi_{8}$ $\neg(\mathsf{f}\lor\mathsf{g})\ \;\mathsf{U}\;\ \left(\mathsf{a}\land\left(\neg\mathsf{b}\;\mathsf{U}\;\mathsf{c}\right)\land\mathsf{F}\,\left(\mathsf{d}\land\left(\neg\mathsf{e}\;\mathsf{U}\;\mathsf{f}\right)\right)\right)$ $\varphi_{16}$ $\mathsf{F}\,\left(\mathsf{b}\land\left(\neg\mathsf{y}\;\mathsf{U}\;(\mathsf{g}\land\mathsf{F}\,(\mathsf{y}\land(\neg(\mathsf{m}\lor\mathsf{g})\;\mathsf{U}\;\mathsf{b})))\right)\right)$
Infinite-horizon $\psi_{1}$ $\mathsf{G}\,\mathsf{F}\,(\mathsf{e}\land(\neg\mathsf{a}\ \;\mathsf{U}\;\ \mathsf{f}))\land\mathsf{G}\,\neg(\mathsf{c}\lor\mathsf{d})$ $\psi_{4}$ $\mathsf{G}\,\mathsf{F}\,\mathsf{b}\land\mathsf{G}\,\mathsf{F}\,\mathsf{g}\land\mathsf{G}\,\neg(\mathsf{y}\lor\mathsf{m})$
$\psi_{2}$ $\mathsf{G}\,\mathsf{F}\,\mathsf{a}\land\mathsf{G}\,\mathsf{F}\,\mathsf{b}\land\mathsf{G}\,\mathsf{F}\,\mathsf{c}\land\mathsf{G}\,\neg(\mathsf{e}\lor\mathsf{f}\lor\mathsf{i})$ $\psi_{5}$ $\mathsf{G}\,\mathsf{F}\,\mathsf{b}\land\mathsf{G}\,\mathsf{F}\,\mathsf{y}\land\mathsf{G}\,\mathsf{F}\,\mathsf{g}\land\mathsf{G}\,\neg\mathsf{m}$
$\psi_{3}$ $\mathsf{G}\,\mathsf{F}\,\mathsf{c}\land\mathsf{G}\,\mathsf{F}\,\mathsf{a}\land\mathsf{G}\,\mathsf{F}\,(\mathsf{e}\land(\neg\mathsf{f}\ \;\mathsf{U}\;\ \mathsf{g}))\land\mathsf{G}\,\mathsf{F}\,\mathsf{k}\land\mathsf{G}\,\neg(\mathsf{i}\lor\mathsf{j})$ $\psi_{6}$ $\mathsf{F}\,\mathsf{G}\,\mathsf{y}\land\mathsf{G}\,\neg(\mathsf{g}\lor\mathsf{b}\lor\mathsf{m})$